OpenStudy (trixitang):

Box A has a volume of 32 cubic meters. Box B is similar to Box A. To create Box B, Box A's dimensions were tripled. What is the volume of Box B? (1 point) 864 m3 288 m3 96 m3 32 m3

OpenStudy (trixitang):

@sammixboo @Er.Mohd.AMIR @paki

OpenStudy (er.mohd.amir):

vol of cube =a*a*a where a is side vol of A =32 find side that is a now box B has side b=3a vol B= b*b*b

OpenStudy (trixitang):

i thinkk i already did it wrong i multiplied 32 three times

OpenStudy (er.mohd.amir):

vol of B = b*b*b=3a*3a*3a use this

OpenStudy (robtobey2):

32*3^3=864